• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Markowitz

    Log entry time 00:00:03 on July19,2003

    Entry number 104613

    keyword=swing shift summary

    started shift in controlled access for investigation of dump ion chambers

    16:09 MCC calls to say beam is on the way. Empty target is currently installed.
    16:13 Harp scan 2153 taken. Horizontal sigma was 451 and vertical was 90. RC Bodo is not satisfied with either number (should be between 100 and 200 for each).

    16:30 Harp scan 2153 shows horizontal ws 490 and vertical was 17
    16:39 Harp scan 2154 shows horizontal sigma was 469 and vertical was 168
    16:46 Harp scan 2156 shows horizontal sigma was 463 and vertical was 32
    16:50 Harp scan 2157 shows horizontal sigma was 407 and vertical was 124
    16:55 Harp scan 2159 shows horizontal sigma was 600 and vertical was 137
    (many adjustments and harp scans later)
    17:19 beam is centered scan 2166, horizontal sigma is 212 and vertical is 109

    Setting up for helicity measurements using Moller, Compton, HRSL, HRSR and Parity DAQs simultaneously.

    18:33 Helicity measurements done. Parity and right HRS agree well, others not
    reported yet.

    Halfwave plate is out, reference target is in with 140 psig, 2 uA
    18:42 start runs 2419(right HRS) and 22415(left HRS). Raster is 2.5(h) by 3(v). Ask MCC to reduce vertical size but keep running.

    18:50 MCC lost a zone and will be down for a few minutes.
    18:58 start runs 2420 (HRSR) and 22416 (HRSL) and 3322(parity). Raster is 2.5(h)
    by 2.5(v). Nitrogen reference cell at 140 psig, 2 uA.
    runs through 2424 are the same thing.
    19:15 pumping the reference cell and doing EPR (36.6%) /NMR (38%) on target

    19:45 starting 3He elastics. half-wave plate is out, target at 0 degrees. Starting runs
    2421 and 22417 and 3323 (right, left, parity DAQ)
    20:05 start runs 2422, 22418, and 3324 (right, left, parity). Still half-wave plate
    out, target at 0 degrees.
    20:32 start runs 2423, 22419, 3325. half-wave plate out, target 0 degrees. runs 2425 same thing

    21:35 ask mcc to insert half-wave plate.

    21:38 start run 2426, 22422, 3328 half-wave plate in, target 0 degrees.
    runs 2426-2432 are all half-wave plate in, target 0 degrees.

    23:27 Moving to reference cell.
    23:33 Empty reference cell, run 2433 (high prescales) and 2434 (low prescales, 1 1)
    23:48 fill reference cell to ~34-35 psig 3He. Set ps 5,1 and start run 2435

    23:59 NO BEAM. MCC says it lost a power supply in the third arc.